OpenStudy (anonymous):

The width of a rectangle is half the length. The perimeter of the rectangle is 54 in. What is the length of the rectangle?

OpenStudy (danjs):

Width = 1/2 * Length Perimeter = 2*Length + 2* Width

OpenStudy (danjs):

They give you the perimeter, sub wideth = 1/2*Length into the perimeter equation. Solve for the Length

OpenStudy (danjs):

Perimeter = 2*Length + 2*(1/2)*Length perimeter is given, find length

OpenStudy (anonymous):

lets call the width \(W\) so the length is \(2W\) the perimeter is \[W+W+2W+2W=6W\] which you know is \(54\) solve \[6W=54\]